A bicycle wheel of radius 50.0cm is spinning at a constant angular speed of 140 rpm in a vertical plane. Find its angular speed.
1
Expert's answer
2020-11-27T10:07:57-0500

The angular speed is


"\omega=140\ rpm=140\frac{2\pi}{60}=14.66\frac{rad}{s}"